# Multiview Equivariance Improves 3D Correspondence Understanding with Minimal Feature Finetuning

### v1

@@ Abstract
Vision foundation models, particularly the ViT family, have revolutionized image understanding by providing rich semantic features. However, despite their success in 2D comprehension, their abilities on grasping 3D spatial relationships are still unclear. In this work, we evaluate and enhance the 3D awareness of ViT-based models. We begin by systematically assessing their ability to learn 3D equivariant features, specifically examining the consistency of semantic embeddings across different viewpoints. Our findings indicate that improved 3D equivariance leads to better performance on various downstream tasks, including pose estimation, tracking, and semantic transfer. Building on this insight, we propose a simple yet effective finetuning strategy based on 3D correspondences, which significantly enhances the 3D correspondence understanding of existing vision models. Remarkably, even finetuning on a single object for just one iteration results in substantial performance gains. All code and resources will be made publicly available to support further advancements in 3D-aware vision models. Our code is available at https://github.com/qq456cvb/3DCorrEnhance.
@@ 2.1.1 Task Definitions
We evaluate on the OnePose-LowTexture and YCB-Video datasets. OnePose-LowTexture He et al. (2022b) includes 40 low-textured household items captured in two videos: one for reference and one for testing, simulating a one-shot scenario. Following He et al. (2022b), pose accuracy is evaluated using 1cm-1deg, 3cm-3deg, and 5cm-5deg thresholds. The YCB-Video dataset Xiang et al. (2017) comprises 21 objects and 92 RGB-D video sequences with pose annotations and CAD models for one-shot generalization. A database is created by rendering objects from 96 icospherical viewpoints. We report Average Recall (AR) for Visible Surface Discrepancy (VSD), Maximum Symmetry-Aware Surface Distance (MSSD), and Maximum Symmetry-Aware Projection Distance (MSPD) following Hodaň et al. (2020).
@@ 2.1.2 On the Choice of Three Tasks
Correspondence estimation is a fundamental component of 3D vision understanding, underlying key tasks such as epipolar geometry, stereo vision for 3D reconstruction, and optical flow or tracking to describe the motion of a perceived 3D world. Stereo cameras, and even human perception, rely on disparity maps—effectively, correspondences between projected 3D parts to understand depth and spatial relationships. The three tasks we evaluated—pose estimation, video tracking, and semantic correspondence—are intentionally selected to cover diverse aspects of correspondence estimation, ranging from simpler to more complex scenarios: 1. Pose Estimation examines correspondences within the same instance under rigid transformations ($SE(3)$); 2. Video Tracking extends this to correspondences for the same instance under potential non-rigid or articulated transformations, such as humans or animals in motion; 3. Semantic Correspondence requires correspondences across different instances with similar semantics, often under arbitrary viewpoint changes. An qualitative illustration of these correspondence types is shown in Figure 4.
@@ 3 Feature Finetuning with Multiview Equivariance
In terms of architecture, besides the common practice of using LoRA to finetune large foundation models, we introduced a single convolutional layer with a kernel size of 3 and a stride of 1. The motivation behind this addition is rooted in the observation that ViT-family models process image tokens as patches, resulting in much lower-resolution feature maps (e.g., 14x smaller in DINOv2). The standard approach to obtain high-resolution per-pixel features is to apply linear interpolation. Consequently, it is beneficial to explicitly exchange information between neighboring patches before interpolation to achieve more accurate results. More ablation results are given in Section 4.1.
@@ 7 Acknowledgements
Yang You, Congyue Deng and Leonidas Guibas acknowledge support from the Toyota Research Institute University 2.0 Program, ARL grant W911NF-21-2-0104, a Vannevar Bush Faculty Fellowship, and a gift from the Flexiv corporation. Yue Wang acknowledges funding supports from Toyota Research Institute, Dolby, and Google DeepMind. Yue Wang is also supported by a Powell Research Award.
@@ A.4 Quantitative Results for Other Variants of DINOv2
OnePose-LowTex TAP-VID-DAVIS PF-PASCAL
1cm 1deg 3cm 3deg 5cm 5deg AJ $\delta_{\text{avg}}$ OA PCK0.05 PCK0.10 PCK0.15
@@ A.8.1 Ablations on Multi-Layer Feature Fusion
In addition to extracting features solely from the last layer, we experiment with two different variations: concatenating the features from the last 4 layers and concatenating features from the 2nd, 5th, 8th, and 11th layers. The results are presented in the Table. We find that fusing features from different layers does improve the instance-level correspondence a little bit but greatly harms semantic correspondences in tracking and semantic transfer. This indicates that features from earlier layers focus more on instance-level details, while the final layer captures more semantic information.
@@ A.9 Other Findings and Discussions
Unlike a perfect sphere, the hemisphere we used is not completely symmetric and provides information about edges and viewpoint orientation. Our visualization of the learned embeddings in Figure 16 shows that after fine-tuning on the hemisphere, the network achieves better edge correspondences and can differentiate between inward and outward views. Even though the object lacks texture, the shadows and edge features provide sufficient cues for the ViT features to develop 3D understanding. Similarly, in cognitive science, scientists have discovered that the human brain excels at inferring 3D structure. Biederman’s Recognition-by-Components (RBC) theory Biederman (1987) suggests that humans recognize objects through simple 3D primitives called geons (geometrical ions)—basic shapes such as cubes, cylinders, and cones.

### v2

@@ Abstract
Vision foundation models, particularly the ViT family, have revolutionized image understanding by providing rich semantic features. However, despite their success in 2D comprehension, their abilities on grasping 3D spatial relationships are still unclear. In this work, we evaluate and enhance the 3D awareness of ViT-based models. We begin by systematically assessing their ability to learn 3D equivariant features, specifically examining the consistency of semantic embeddings across different viewpoints. Our findings indicate that improved 3D equivariance leads to better performance on various downstream tasks, including pose estimation, tracking, and semantic transfer. Building on this insight, we propose a simple yet effective finetuning strategy based on 3D correspondences, which significantly enhances the 3D correspondence understanding of existing vision models. Remarkably, finetuning on a single object for one iteration results in substantial gains.
@@ 2.1.1 Task Definitions
We evaluate on the OnePose-LowTexture and YCB-Video datasets. OnePose-LowTexture He et al. (2022b) includes 40 low-textured household items captured in two videos: one for reference and one for testing, simulating a one-shot scenario. We evaluate on every 10 frames in the video. Following He et al. (2022b), pose accuracy is evaluated using 1cm-1deg, 3cm-3deg, and 5cm-5deg thresholds. The YCB-Video dataset Xiang et al. (2017) comprises 21 objects and 92 RGB-D video sequences with pose annotations and CAD models for one-shot generalization. A database is created by rendering objects from 96 icospherical viewpoints. We report Average Recall (AR) for Visible Surface Discrepancy (VSD), Maximum Symmetry-Aware Surface Distance (MSSD), and Maximum Symmetry-Aware Projection Distance (MSPD) following Hodaň et al. (2020).
@@ 2.1.2 On the Choice of Three Tasks
Correspondence estimation is a fundamental component of 3D vision understanding, underlying key tasks such as epipolar geometry, stereo vision for 3D reconstruction, and optical flow or tracking to describe the motion of a perceived 3D world. Stereo cameras, and even human perception, rely on disparity maps—effectively, correspondences between projected 3D parts to understand depth and spatial relationships.
The three tasks we evaluated—pose estimation, video tracking, and semantic correspondence—are intentionally selected to cover diverse aspects of correspondence estimation, ranging from simpler to more complex scenarios: 1. Pose Estimation examines correspondences within the same instance under rigid transformations ($SE(3)$); 2. Video Tracking extends this to correspondences for the same instance under potential non-rigid or articulated transformations, such as humans or animals in motion; 3. Semantic Correspondence requires correspondences across different instances with similar semantics, often under arbitrary viewpoint changes. An qualitative illustration of these correspondence types is shown in Figure 4.
@@ 3 Feature Finetuning with Multiview Equivariance
In terms of architecture, we apply LoRA in the last four blocks to finetune large foundation models, we introduced a single convolutional layer with a kernel size of 3 and a stride of 1. The motivation behind this addition is rooted in the observation that ViT-family models process image tokens as patches, resulting in much lower-resolution feature maps (e.g., 14x smaller in DINOv2). The standard approach to obtain high-resolution per-pixel features is to apply linear interpolation. Consequently, it is beneficial to explicitly exchange information between neighboring patches before interpolation to achieve more accurate results. More ablation results are given in Section 4.1.
@@ 6 Conclusion
7 Statements
Ethics Statement. Our method leverages open-sourced simulation data and real data whose data collection process follows strict ethical guidelines. In using these data, we follow the same ethical considerations to protect sensitive information. There is no ethical concerns detected of the proposed method to our knowledge, and we will strive to adhere to ICLR code of conduct for future use of the proposed method.
Reproducibility Statement. We provide extensive details for ease of re-implementation. We strive to ensure our method is reproducible and the findings in this paper are generalizalble. We will release code, results, and scripts for reproduction to promote future research in 3D deep learning.
@@ 8 Acknowledgements
Yang You is supported in part by the Outstanding Doctoral Graduates Development Scholarship of Shanghai Jiao Tong University. Congyue Deng, Yang You, and Leonidas Guibas acknowledge support from the Toyota Research Institute University 2.0 Program, ARL grant W911NF-21-2-0104, a Vannevar Bush Faculty Fellowship, and a gift from the Flexiv corporation. Yue Wang acknowledges funding supports from Toyota Research Institute, Dolby, and Google DeepMind. Yue Wang is also supported by a Powell Research Award.
@@ A.4 Quantitative Results for Other Variants of DINOv2
OnePose-LowTex TAP-VID-DAVIS PF-PASCAL (Diff. View)
1cm-1deg 3cm-3deg 5cm-5deg AJ $\delta_{\text{avg}}$ OA PCK0.05 PCK0.10 PCK0.15
@@ A.8.1 Ablations on Multi-Layer Feature Fusion
In addition to extracting features solely from the last layer (11th), we experiment with two different variations: concatenating the features from the last 4 layers and concatenating features from the 2nd, 5th, 8th, and 11th layers. The results are presented in the Table 16. We find that fusing features from different layers does improve the instance-level correspondence a little bit but greatly harms semantic correspondences in tracking and semantic transfer. This indicates that features from earlier layers focus more on instance-level details, while the final layer captures more semantic information.
OnePose-LowTex TAP-VID-DAVIS PF-PASCAL (Diff. View)
1cm-1deg 3cm-3deg 5cm-5deg AJ $\delta_{\text{avg}}$ OA PCK0.05 PCK0.10 PCK0.15
Layer 11 13.58 58.03 77.35 46.85 63.84 84.15 47.24 60.76 67.57
Layer 2,5,8,11 15.34 59.56 76.81 39.67 56.74 76.29 39.84 53.05 60.15
Layer 8,9,10,11 14.24 60.35 79.27 41.25 56.56 80.15 44.99 57.73 64.48
Table 16: Ablations on the choice of feature blocks on DINOv2 base model.
@@ A.9 Other Findings and Discussions
Unlike a perfect sphere, the hemisphere we used is not completely symmetric and provides information about edges and viewpoint orientation. Our visualization of the learned embeddings in Figure 16 shows that after fine-tuning on the hemisphere, the network achieves better edge correspondences and can differentiate between inward and outward views. Even though the object lacks texture, the shadows and edge features provide sufficient cues for the ViT features to develop 3D understanding.
Similarly, in cognitive science, scientists have discovered that the human brain excels at inferring 3D structure. Biederman’s Recognition-by-Components (RBC) theory Biederman (1987) suggests that humans recognize objects through simple 3D primitives called geons (geometrical ions)—basic shapes such as cubes, cylinders, and cones.